OpenStudy (anonymous):

How to find t? (4t + 3)^(-1/2) = -3 How do you do this if the exponent is a negative fraction?

OpenStudy (agent0smith):

a^(-1) = 1/a

OpenStudy (anonymous):

Yes, I know that...

OpenStudy (agent0smith):

Or, just put both sides to the power of -2.

OpenStudy (anonymous):

If you do that, is the left side left with 4t+3?

OpenStudy (agent0smith):

Yes, right will be (-3)^(-2)

OpenStudy (anonymous):

You can take -3 as -1(3)^1/2

OpenStudy (anonymous):

That's so confusing...

OpenStudy (anonymous):

If i do what agent0smith said to do, is it correct if i get 4t+3 = -1/9 ?

OpenStudy (agent0smith):

^no, (-3)^2 = 9 not -9

OpenStudy (anonymous):

Sorry You can take -3 as -1(9)^1/2

OpenStudy (anonymous):

Okay. I will solve from 4t+3 = 1/9 then! :)

OpenStudy (mathstudent55):

@esl.0095 the right side is (-3)^(-2) = 1/9

OpenStudy (anonymous):

Yes, I got that :) thank you :)

OpenStudy (anonymous):

So final answer... t = -13/18. Anyone interested?

OpenStudy (agent0smith):

yep

OpenStudy (anonymous):

Thanks agent0smith and all of you others. I will close the quesiton now! :D

OpenStudy (mathstudent55):

The solution is not \(t = -\dfrac{13}{18} \) Let's look at the first few steps of dealing with a negative exponent and a fractional exponent. \(\large (4t + 3)^{-\frac{1}{2}} = -3\) \(\large \left(\dfrac{1}{4t + 3} \right)^{\frac{1}{2}} = -3\) \(\large \sqrt{\dfrac{1}{4t + 3}} = -3\) Do you understand the steps so far?

OpenStudy (agent0smith):

Oh yeah, you may have to check your solution since we squared both sides

OpenStudy (mathstudent55):

Notice that you have a square root equaling a negative number. Therefore, there is no real solution.

OpenStudy (agent0smith):

Yep. Btw @mathstudent55 nothing you wrote shows up.

OpenStudy (mathstudent55):

Correct. When you raise both sides to the -2 power you are both squaring and getting the reciprocal. Any time you square both sides of an equation, you need to check the solution in the original equation because squaring both sides of an equation may introduce extraneous solutions.

OpenStudy (mathstudent55):

@agent0smith I see what I wrote in my previous responses, and LaTeX is showing properly on my screen.

OpenStudy (agent0smith):

What browser are you on? Doesn't work on chrome

OpenStudy (anonymous):

Guys, I made a mistake prior to this, so I didn't need to solve for (4t + 3)^(-1/2) = -3 in the first place. Sorry and thanks to all you still for helping me!! We can close this question for good now :)